OpenStudy (anonymous):

graph x^2-y^2=16 what are its lines of symmetry?

OpenStudy (anonymous):

@ranga i know you got me

OpenStudy (ranga):

Write the equation in standard form: \[\frac{ x ^{2} }{ 16 } - \frac{ y ^{2} }{ 16} = 1\] \[\frac{ x ^{2} }{ 4^{2} } - \frac{ y ^{2} }{ 4^{2}} = 1\]  Hard to draw hyperbolas without the proper drawing tool. Anyway, I think the lines of symmetry are both the x axis and the y axis.